OpenStudy (anonymous):

Use the function to answer the question. h : k -> k^2 - k What is the value of h(1)? A) -1 B) 0 C) 1 D) 2

OpenStudy (raden):

it means that function is h(k)=k^2 - k so, just put k=1 to get the value of h(1)

OpenStudy (raden):

h(1) = 1^2 - 1 = ... ?

OpenStudy (anonymous):

is it 0?

OpenStudy (raden):

yes, correct

OpenStudy (anonymous):

ty mate :)

OpenStudy (raden):

welcome, friend :)